A professor decided to create a student committee of three students. Initially he selects 5 females and 4 males. How many possible committees can be created with 3 males and how many can be created with 2 females?
Answer by edjones(8007)

Each committee will have 3 people on it.
3 males: 4C3=4
2 females + 1 male: 5C2 * 4C1 = 10*4 = 40
.
Ed